The consumer price index (CPI) rose from 106.7 to 124.9, while Helen's net salary didn't change at all. Calculate the relative real change in Helen's salary (purchasing power) as percentage.

1
Expert's answer
2022-02-14T16:23:52-0500

Relative change is calculated by subtracting the value of the indicator in the first period from the value of the indicator in the second period which is then divided by the value of the indicator in the first ..


124.9-106.7=18 .2

18.2/106.7=0.17057

0.17057×100=17.057%